### v1

# Newtonian potential from scattering amplitudes in super-renormalizable gravity

Abstract
Using scattering amplitudes, we compute the coupling between a general super-renormalizable gravity theory and massive scalar particles. This allows us to derive the $D$-dimensional Newtonian potential at both tree-level and one-loop level—the latter containing the first calculation by using newly derived three-graviton Feynman rules. In four-dimensional spacetime, we numerically demonstrate that the Newtonian potential remains finite at the origin, providing compelling evidence that the singularity-free nature of super-renormalizable gravity persists at the one-loop level.
Introduction. The quest to reconcile quantum mechanics with general relativity (GR) is not merely a technical issue of taming ultraviolet (UV) divergences, but must also address the fate of the classical singularities. A powerful bridge to this quantum-classical interface is provided by the effective field theory approach, where quantum field theory (QFT) techniques are deployed to extract classical gravitational dynamics from quantum amplitudes Duff:1973zz ; Bjerrum-Bohr:2018xdl ; Goldberger:2004jt ; Kosower:2018adc ; Cheung:2018wkq ; Bern:2019nnu ; Chung:2019yfs ; Donoghue:1993eb ; Bjerrum-Bohr:2002fji ; Donoghue:1994dn ; Donoghue:1993eb ; Bjerrum-Bohr:2002fji ; Neill:2013wsa ; Damour:2016gwp . In this framework, GR emerges from the exchange of spin-2 gravitons Donoghue:1994dn , and modern on-shell methods, such as scattering amplitudes and generalized unitarity, have dramatically refined our ability to reveal the deeper connections between QFT and spacetime dynamics Bern:2008qj ; Mangano:1990by ; Aoude:2023fdm ; DiVecchia:2021bdo ; Brandhuber:2021eyq ; Monteiro:2021ztt ; Manohar:2022dea ; Aoude:2022trd ; Menezes:2022tcs ; Kalin:2022hph ; Guevara:2019fsj ; Aoude:2021oqj ; Herrmann:2021lqe ; Cristofoli:2020uzm .
Scattering amplitudes provide a uniquely powerful framework for probing quantum gravity (QG), with gauge-invariant observables laying the foundation for a future QG phenomenology Kalin:2020fhe ; Bern:2021yeh ; Bern:2021dqo ; Adamo:2022qci ; Bohnenblust:2024hkw ; Jakobsen:2021lvp ; Mougiakakos:2024nku . Their theoretical precision enables the systematic extraction of QG effects, including the constraining of higher-derivative operators induced by loop divergences in effective approaches Wang:2015jna . While tree-level amplitudes already encode rich physical content, their direct computation is often obscured by gauge redundancies and field-redefinition ambiguities in the off-shell Lagrangian formulation DeWitt:1967uc ; Berends:1974gk . To overcome these challenges, modern amplitude methods-most notably generalized unitarity-have been developed, which isolate the physical, “cut-constructible” parts of loop amplitudes by focusing on their non-analytic structure Bern:1994zx . This approach has proven transformative, enabling efficient calculations of long-range interactions in gravity that are governed precisely by such non-analytic terms, scaling as $m/\sqrt{-q^{2}}$ for a massive probe $m$ and momentum transfer $q$ Donoghue:1993eb . These technical advances now drive concrete progress in classical GR, facilitating high-precision computations of post-Newtonian and post-Minkowskian dynamics in binary systems Blanchet:2006jqj .
Super-renormalizable gravity provides a consistent and predictive framework for ultraviolet-complete QG, preserving unitarity, covariance, and perturbative control across all scales Biswas:2011ar ; Modesto:2011kw ; Briscese:2013lna ; Tomboulis:2015esa ; Calcagni:2014vxa ; Alexander:2012aw ; Dona:2015tra . Its defining feature is weak nonlocality—a property shared by string theory Brekke:1987ptq —which ensures super-renormalizability while excluding pathological states. Although traditional perturbative studies in super-renormalizable gravity have established a finite Newtonian potential at $\mathcal{O}(G)$ order Biswas:2011ar and explored other physical aspects Briscese:2013lna ; Krasnikov:1987yj ; Modesto:2011kw ; Briscese:2013lna ; Tomboulis:2015esa ; Dona:2015tra , the behavior at $\mathcal{O}(G^{2})$ order has remained a critical open question, we resolve this outstanding challenge by applying modern amplitude techniques to perform the first complete one-loop computation of the Newtonian potential.
We focus here on a general super-renormalizable gravity scenario described by the action Krasnikov:1987yj ; Biswas:2011ar ; Modesto:2011kw ; Briscese:2013lna ; Tomboulis:2015esa ; Calcagni:2014vxa ; Alexander:2012aw ; Dona:2015tra
$S=\frac{2}{\kappa^{2}}\int\sqrt{-g}\,\mathrm{d}^{D}x\left(R+RF_{1}(\Box)R+R_{\mu\nu}F_{2}(\Box)R^{\mu\nu}+R_{\mu\nu\alpha\beta}F_{4}(\Box)R^{\mu\nu\alpha\beta}\right),$ (1)
where $R$ denotes the Ricci scalar curvature and $\kappa^{2}=32\pi G_{N}$. The functions of the covariant d’Alembertian operator $F_{i}(\Box)$, called form factors, are assumed to be entire functions and can therefore be expanded in a Taylor series as $F_{i}(\Box)=\sum^{\infty}_{n=0}f_{in}\frac{\Box^{n}}{M^{2n}_{\ast}}$ ($i=1,2$), where $M_{\ast}$ denotes the mass scale at which the higher-derivative terms in the action become relevant. The form factors $F_{i}(\Box)$ can be expressed in terms of exponentials of entire functions $H_{\ell}(\Box)$ (with $\ell=0,2$) Koshelev:2017ebj .
$\begin{split}&F_{1}(\Box)=-\frac{(D-2)\left(e^{H_{0}(\Box)}-1\right)+D\left(e^{H_{2}(\Box)}-1\right)}{4(D-1)\Box}+F_{4}(\Box),\\ &F_{2}(\Box)=\frac{e^{H_{2}(\Box)}-1}{\Box}-4F_{4}(\Box).\end{split}$ (2)
The requirement of super-renormalizability imposes two key constraints. First, the form factor $F_{4}(\Box)$ must exhibit the same UV asymptotic behavior as $F_{i}(\Box)$ ($i=0,2$), a condition most straightforwardly satisfied by setting $F_{4}(\Box)=0$, thereby reducing to two independent form factors. Second, the polynomials defining $H_{0}(\Box)$ and $H_{2}(\Box)$ must share the same degree, while the entire functions $e^{H_{i}(\Box)}$ must simultaneously fulfill three distinct consistency conditions Tomboulis:2015esa .
Newtonian potential at tree-level. The classical limit of QFT, corresponding to $\hbar\to 0$, is conventionally identified with the stationary-point approximation $\delta S=0$. Nevertheless, Feynman diagrams to all loop orders can still contribute classically, as comprehensively discussed in Kosower:2018adc and specifically in the gravitational context Bjerrum-Bohr:2018xdl ; Cheung:2020gyp ; Bjerrum-Bohr:2019kec . In this framework, we analyze the vertex function for massive scalars interacting with gravitons: the scalars are treated as point particles with finite momenta, while gravitons carry soft momenta that vanish in the classical limit. This scaling applies to both external and loop momenta, implying that the classical limit corresponds physically to the long-range limit of the interaction. The vertex rules come from the expansion of the matter term of the action which is
$S_{\text{matter}}=\frac{1}{2}\int\,\mathrm{d}^{D}x\sqrt{-g}\left(\nabla_{\mu}\phi\nabla^{\mu}\phi-m^{2}\phi^{2}\right),$ (3)
with tree-level Feynman diagram
Figure 1: The tree-level contribution to the three-point vertex function involves the graviton momentum $q$, and the four-momentum $k$ of the on-shell particle, which satisfies the condition $k^{2}=m^{2}$.
In calculations, it is convenience to separate tensors into parallel and orthogonal components relative to the momentum $k_{\mu}$ of the scalar particle by using the projection operators: $\eta^{||}_{\mu\nu}=\frac{k_{\mu}k_{\nu}}{m^{2}}$, and $\eta^{\bot}_{\mu\nu}=n_{\mu\nu}-\frac{k_{\mu}k_{\nu}}{m^{2}}$, which are diagonal in the inertial frame of $k_{\mu}$ and represent the time and spatial components of $\eta_{m\nu}$, respectively.
In the classical limit, contributions from Feynman diagrams persist to all loop orders, with each additional loop introducing higher powers of the gravitational coupling $G_{N}$. Physically, the long-range interaction is dominated by the region where graviton momentum $q\to 0$, corresponding to a heavy scalar source of mass $m$ at rest. As depicted in Fig. 1, the amplitude in this regime encodes the metric perturbation $h_{\mu\nu}$ generated by the point-like source and its gravitational field. This amplitude, denoted $\mathcal{M}^{\mu\nu}_{\text{vertex}}$, incorporates both the energy-momentum pseudotensor and gauge-fixing contributions. A relation derived in Jakobsen:2020ksu links the gauge-invariant part of the scattering amplitude to the physical pseudotensor as
$2\pi\delta\left(kq\right)\mathcal{M}^{\mu\nu}_{\text{vertex}}=-\kappa\tilde{\tau}^{\mu\nu}+\frac{1}{\xi}\tilde{H}^{\mu\nu}_{\text{non-linear}},$ (4)
where $\tilde{\tau}^{\mu\nu}$ represents the combined contribution from the matter ($T_{\mu\nu}$) and the nonlinear gravitational field. To zeroth order, it corresponds to the energy-momentum tensor of point particle in special relativity, while loop corrections account for the energy-momentum contribution from the surrounding, self-interacting gravitational field. The term $\tilde{H}^{\mu\nu}_{\text{non-linear}}$ arises from the nonlinear part of the gauge-fixing term is chosen as: $S_{\text{gf}}=\int\mathrm{d}^{D}x\,\eta^{\mu\nu}\partial_{\alpha}h^{\alpha}_{\mu}\,\omega(\Box)\,\partial_{\beta}h^{\beta}_{\nu}/(\kappa^{2}\xi)$. Due to the conservation of momentum in both the initial and final states, the constraint $\delta(kq)=\delta(mq_{||})$ must be imposed to properly define the integral in classical limit. To obtain the correction to the metric, we perform a Fourier transform of the above expression and combine it with the equations of motion, resulting in
$g_{\mu\nu}=\eta_{\mu\nu}-\frac{\kappa}{2}\int\frac{\,\mathrm{d}^{D}q\delta(kq)e^{-iqx}}{(2\pi)^{D-1}}O^{-1}_{\alpha\beta\mu\nu}\mathcal{M}_{\mathrm{vertex}}^{\alpha\beta},$ (5)
where $O^{-1}_{\alpha\beta\mu\nu}$ represents the propagator in gravitational theory. For GR, it takes the form $O^{-1}_{\alpha\beta\mu\nu}=\frac{I_{\alpha\beta\mu\nu}-\frac{1}{D-2}\eta_{\mu\nu}\eta_{\alpha\beta}}{q^{2}+i\epsilon}$, corresponding to the Schwarzschild-Tangherlini solution Jakobsen:2020ksu . For the sake of simplifying calculations, we examine a particular class of form factors: $\omega(\Box)=e^{-\Box/M^{2}_{\ast}}$, and $H_{0}(\Box)=H_{2}(\Box)=-\Box/M^{2}_{\ast}$. Therefore $F_{1}(\Box)=-F_{2}(\Box)/2=\left(e^{-\Box/M^{2}_{\ast}}-1\right)/\Box$. It is straightforward to show that these form factors simplify the propagator as
$\begin{split}O^{-1}_{\mu\nu\alpha\beta}&=\frac{e^{-\frac{q^{2}}{M^{2}_{\ast}}}}{q^{2}}I_{\mu\nu\alpha\beta}-\frac{4e^{-\frac{q^{2}}{M^{2}_{\ast}}}}{(D-2)q^{2}}T_{\mu\nu\alpha\beta}+\frac{2e^{-\frac{q^{2}}{M^{2}_{\ast}}}}{(D-2)q^{2}}C_{\mu\nu\alpha\beta}\\ &+\frac{(2\xi-2)e^{-\frac{q^{2}}{M^{2}_{\ast}}}}{q^{2}}J_{\mu\nu\alpha\beta}+\frac{e^{-\frac{q^{2}}{M^{2}_{\ast}}}}{q^{2}}\left(\frac{D-2}{D-1}-\frac{1}{(D-2)(D-1)}-\frac{3\xi}{2}\right)K_{\mu\nu\alpha\beta},\end{split}$ (6)
where the five independent operators are $I_{\alpha\beta}^{\mu\nu}=\frac{1}{2}\left(\delta_{\alpha}^{\mu}\delta_{\beta}^{\nu}+\delta_{\beta}^{\mu}\delta_{\alpha}^{\nu}\right)$, $T_{\alpha\beta}^{\mu\nu}\equiv\frac{1}{4}\eta^{\mu\nu}\eta_{\alpha\beta}$, $C_{\alpha\beta}^{\mu\nu}\equiv\frac{1}{2}\left(\eta^{\mu\nu}\frac{q_{\alpha}q_{\beta}}{q^{2}}+\frac{q^{\mu}q^{\nu}}{q^{2}}\eta_{\alpha\beta}\right)$, $J_{\alpha\beta}^{\mu\nu}\equiv I_{\rho\kappa}^{\mu\nu}\frac{q_{\sigma}q^{\rho}}{q^{2}}I_{\alpha\beta}^{\sigma\kappa}$, and $K_{\alpha\beta}^{\mu\nu}\equiv\frac{q^{\mu}q^{\nu}}{q^{2}}\frac{q_{\alpha}q_{\beta}}{q^{2}}$. With the Feynman rules, the tree-level scattering amplitude is given by $\mathcal{M}^{\mu\nu}_{\text{tree}}=-\kappa m^{2}\eta^{||}_{\mu\nu}$, reflecting the leading coupling of massive scalars to gravity. Based on the coupling between the propagator and the scattering amplitudes (5), we calculate the first-order correction to the metric tensor as
$\begin{split}h^{(G_{N})}_{\alpha\beta}&=-\frac{(D-3)4\pi G_{N}m}{(D-2)\pi^{\frac{D-1}{2}}r^{D-3}}\left[\Gamma\left(\frac{D-3}{2}\right)-\Gamma\left(\frac{D-3}{2},\frac{1}{4}M^{2}_{\ast}r^{2}\right)\right]\eta^{||}_{\alpha\beta}\\ &+\frac{4\pi G_{N}m}{(D-2)\pi^{\frac{D-1}{2}}r^{D-3}}\left[\Gamma\left(\frac{D-3}{2}\right)-\Gamma\left(\frac{D-3}{2},\frac{1}{4}M^{2}_{\ast}r^{2}\right)\right]\eta^{\bot}_{\alpha\beta},\end{split}$ (7)
which gives the $D$-dimensional Newtonian potential as
$\begin{split}\Phi^{(G_{N})}(r)=-\frac{(D-3)2\pi G_{N}m}{(D-2)\pi^{\frac{D-1}{2}}r^{D-3}}\left[\Gamma\left(\frac{D-3}{2}\right)-\Gamma\left(\frac{D-3}{2},\frac{1}{4}M^{2}_{\ast}r^{2}\right)\right].\end{split}$ (8)
In particular, for $D=4$, the Newtonian potential reduces to $\Phi^{(G_{N})}(r)=-\frac{G_{N}m}{r}\text{erf}\left(\frac{M_{\ast}r}{2}\right)$, where the error function is defined as $\text{erf}(x)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}e^{-t^{2}}dt$. This result aligns with earlier work in Biswas:2011ar obtained with linear perturbation. The error function encodes the nonlocal smearing of the gravitational interaction: at large distances $\lim_{x\to\infty}\text{erf}(x)=1$, the potential reduces to the standard Newtonian form $\frac{G_{N}m}{r}$, recovering GR in the infrared. In the short-range regime, however, the potential remains perfectly regular: $\Phi^{(G_{N})}(r\to 0)=-\frac{G_{N}mM_{*}}{\sqrt{\pi}}$. The weak nonlocality regulates the ultraviolet behavior of gravity, defusing the classical singularity at tree level.
Newtonian potential at one-loop level. At order $(G_{N})^{2}$, the classical metric correction originates from the triangle one-loop diagram (Fig. 2), which uniquely provides the non-analytic terms responsible for long-range interactions Bjerrum-Bohr:2018xdl . Other one-loop topologies, devoid of such non-analyticity, do not contribute to the classical potential, thereby establishing the triangle diagram as the sole contributor to the $\mathcal{O}(G_{N}^{2})$ singularity resolution.
Figure 2: The one-loop contribution to the three-point vertex function involves $q$, and $k$. The loop momentum is denoted by $l$.
Employing the Feynman rules in Appendix A for the three-graviton vertex and for the interaction between two scalar particles and one graviton, together with the propagator from Eq. (6) and the metric correction in Eq. (5), we obtain
$\begin{split}i\mathcal{M}^{\mu\nu}_{\text{vertex}}&\equiv i\mathcal{M}^{\mu\nu}_{\text{GR}}+i\mathcal{M}^{\mu\nu}_{\text{Nonlocal}}\\ &=-2m^{4}\kappa^{3}\int\frac{{\,\mathrm{d}^{D}l}}{(2\pi)^{D}}\frac{e^{-\frac{l^{2}}{M^{2}_{\ast}}}e^{-\frac{\left(l+q_{\bot}\right)^{2}}{M^{2}_{\ast}}}f_{\alpha\beta}f_{\gamma\delta}\left(V^{\alpha\beta\gamma\delta\mu\nu}_{h^{3}\text{GR}}+V^{\alpha\beta\gamma\delta\mu\nu}_{h^{3}\text{Nonocal}}\right)}{l^{2}(l+q_{\bot})^{2}\left((l+k)^{2}-m^{2}+i\epsilon\right)},\end{split}$ (9)
where $V^{\alpha\beta\gamma\delta\mu\nu}_{h^{3}\text{GR}}$ and $V^{\alpha\beta\gamma\delta\mu\nu}_{h^{3}\text{Nonlocal}}$ denote the three-graviton vertex interactions originating from the Einstein–Hilbert term and the nonlocal sector of the action, respectively. A key element in the amplitude is the tensor
$f_{\alpha\beta}\equiv\frac{D-3}{D-2}\eta^{\parallel}_{\alpha\beta}-\frac{1}{D-2}\eta^{\perp}_{\alpha\beta},$ (10)
which effectively encodes the scalar–graviton vertices contracted with the graviton propagator. This tensor captures the physical structure inherited from the $\phi^{2}h$ coupling after integrating out intermediate graviton modes.
In the classical limit, where the graviton momenta are soft compared to the momentum of the massive scalar particle, the scalar propagator admits a significant simplification. Specifically, terms involving graviton momenta $l$ and $q$ in the scalar–graviton vertex, as well as the momentum-dependent tensor structures in the propagator numerator, become subdominant and can be systematically neglected. This soft-graviton approximation isolates the leading long-range interaction and is essential for extracting the classical dynamics from the amplitude.
Considering on the tensor part of the integral involving the three-graviton vertex from the GR contribution $\left(V^{\alpha\beta\gamma\delta\mu\nu}_{h^{3}\text{GR}}\right)$ and nonlocal part $\left(V^{\alpha\beta\gamma\delta\mu\nu}_{h^{3}\text{Nonlocal}}\right)$, yields
[…]
where $U$-tensor is defined in Appendix A, and the quantities $I^{\text{GR}}_{a_{1}...a_{n}}$, and $I^{(i)}_{a_{1}...a_{n}}$ ($i=1,2,3$) are given by
$\begin{split}&I^{\text{GR}}_{a_{1}...a_{n}}\equiv\int\frac{{\,\mathrm{d}^{D}l}}{(2\pi)^{D}}\frac{e^{-\frac{l^{2}}{M^{2}_{\ast}}}e^{-\frac{\left(l+q_{\bot}\right)^{2}}{M^{2}_{\ast}}}l_{a_{1}}...l_{a_{n}}}{l^{2}(l+q_{\bot})^{2}\left((l+k)^{2}-m^{2}+i\epsilon\right)},\end{split}$ (12)
$\begin{split}&I^{(1)}_{a_{1}...a_{n}}\equiv-\int\frac{{\,\mathrm{d}^{D}l}}{(2\pi)^{D}}\frac{e^{-\frac{l^{2}}{M^{2}_{\ast}}}e^{-\frac{\left(l+q_{\bot}\right)^{2}}{M^{2}_{\ast}}}\left(e^{\frac{\left(l+q_{\bot}\right)^{2}}{M^{2}_{\ast}}}-1\right)l_{a_{1}}...l_{a_{n}}}{l^{2}(l+q_{\bot})^{4}\left((l+k)^{2}-m^{2}+i\epsilon\right)},\end{split}$ (13)
$\begin{split}&I^{(2)}_{a_{1}...a_{n}}\equiv-\int\frac{{\,\mathrm{d}^{D}l}}{(2\pi)^{D}}\frac{e^{-\frac{l^{2}}{M^{2}_{\ast}}}e^{-\frac{\left(l+q_{\bot}\right)^{2}}{M^{2}_{\ast}}}\left(e^{\frac{q^{2}_{\bot}}{M^{2}_{\ast}}}-1\right)l_{a_{1}}...l_{a_{n}}}{l^{2}q^{2}_{\bot}(l+q_{\bot})^{2}\left((l+k)^{2}-m^{2}+i\epsilon\right)},\end{split}$ (14)
$\begin{split}&I^{(3)}_{a_{1}...a_{n}}\equiv-\int\frac{{\,\mathrm{d}^{D}l}}{(2\pi)^{D}}\frac{e^{-\frac{l^{2}}{M^{2}_{\ast}}}e^{-\frac{\left(l+q_{\bot}\right)^{2}}{M^{2}_{\ast}}}\left(e^{\frac{l^{2}}{M^{2}_{\ast}}}-1\right)l_{a_{1}}...l_{a_{n}}}{l^{4}(l+q_{\bot})^{2}\left((l+k)^{2}-m^{2}+i\epsilon\right)}.\end{split}$ (15)
Ultimately, the corrections to the Newtonian potential up to second order in $(G_{N})^{2}$ within the one-loop approximation can be expressed as
$\begin{split}\Phi^{(G_{N})^{2}}(r)&=-\frac{\kappa^{4}m^{2}}{4}\int\frac{\,\mathrm{d}^{D-1}q_{\bot}}{(2\pi)^{D-1}}e^{-iq_{\bot}x_{\bot}}e^{-\frac{q^{2}_{\bot}}{M^{2}_{\ast}}}\left[N_{D-1}\frac{183D+5D^{3}-61D^{2}-151}{16(D-2)^{2}}\right.\\ &\left.-q^{2}_{\bot}\left(N^{(1)}_{D-1}+N^{(3)}_{D-1}\right)\left(\frac{D^{2}-5D+8}{4(D-2)^{3}}\right)-q^{2}_{\bot}N^{(2)}_{D-1}\left(\frac{76+13D+10D^{2}}{8(D-2)^{3}}\right)\right],\end{split}$ (16)
where the integrals $N_{D-1},N^{(i)}_{D-1},i=1,2,3$ are given in Appendix B. The Newtonian potential at order $\mathcal{O}(G_{N}^{2})$, given in Eq. (16), is structurally governed by the spacetime dimension $D$ and four loop integrals $N_{D-1},N^{(i)}_{D-1}$ (with $i=1,2,3$), whose explicit forms are provided in Appendix B. These integrals do not admit a simple analytic expression, motivating a numerical treatment in the physically relevant case $D=4$. As displayed in Fig. 3, the computed potential is plotted against the dimensionless radial coordinate, the one-loop correction grows smoothly with distance, matching expected long-range behavior. Most notably, the potential remains completely regular as $r\to 0$, offering direct numerical evidence that the mechanism of singularity resolution, enforced by weak nonlocality, persists at the quantum level.
Figure 3: The Newtonian potential in $D=4$ dimension as a function of distance.
Conclusion. We have investigated the Newtonian potential in super-renormalizable gravity through scattering amplitudes, systematically including both tree-level and one-loop contributions in $D$ dimensions. At tree level, the potential in $D=4$ is nonsingular and governed by an error function, thereby regularizing the classical divergence. We further performed the first one-loop calculation of the $\mathcal{O}(G^{2})$ correction, which required deriving new Feynman rules for the three-graviton vertex in the nonlocal framework. Although the one-loop result involves intricate integrals that preclude simple analytic forms, a numerical evaluation in four dimensions confirms unambiguously that the potential remains finite at the origin. This provides the first conclusive evidence that singularity resolution in super-renormalizable gravity holds at the one-loop level. Whether this remarkable property persists at higher loop orders remains an important question for future investigation.
Acknowledgements.
This study was supported by the National Natural Science Foundation of China (Grant No. 12333008).
Appendix A Gravitational Self-Interaction: 3-graviton vertices
We present here the Feynman rules in the covariant harmonic gauge for the three-graviton vertex in super-renormalizable gravity, derived from Eq. (1). We focus on the contributions from the gravitational field alone, gauge-fixing terms are omitted as they do not contribute to the three-point interaction. The vertex is conveniently written in terms of the $U_{\text{GR}}$-tensor introduced in
$\begin{split}S_{\text{GR}}&=\frac{2}{\kappa^{2}}\int\sqrt{-g}\,\mathrm{d}^{D}xR=\kappa\int\,\mathrm{d}^{D}x\tilde{h}_{\mu\nu}U^{\mu\nu\alpha\beta\rho\gamma\delta\sigma}_{\text{GR}}\partial_{\rho}\tilde{h}_{\alpha\beta}\partial_{\sigma}\tilde{h}_{\delta\gamma}\\ &=-\kappa\int\frac{\,\mathrm{d}^{D}l_{1}}{(2\pi)^{D}}\int\frac{\,\mathrm{d}^{D}l_{2}}{(2\pi)^{D}}\int\frac{\,\mathrm{d}^{D}l_{3}}{(2\pi)^{D}}(2\pi)^{D}\delta^{D}\left(l_{1}+l_{2}+l_{3}\right)\tilde{h}_{\mu\nu}(l_{1})U^{\mu\nu\alpha\beta\rho\gamma\delta\sigma}_{\text{GR}}\tilde{h}_{\alpha\beta}(l_{2})l_{2\rho}l_{3\sigma}\tilde{h}_{\delta\gamma}(l_{3}),\end{split}$ (17)
where the tensor $U^{\mu\nu\alpha\beta\rho\gamma\delta\sigma}_{\text{GR}}$ is defined to be symmetric under the exchange of the index groups $\alpha\beta\rho\leftrightarrow\gamma\delta\sigma$, following Fourier transformation in $D$-dimensional spacetime. It further exhibits full symmetry under the individual exchanges $\mu\leftrightarrow\nu$, $\alpha\leftrightarrow\beta$, and $\gamma\leftrightarrow\delta$, and admits the explicit form
$\begin{split}U^{\mu\nu\alpha\beta\rho\gamma\delta\sigma}_{\text{GR}}\tilde{h}_{\mu\nu}\partial_{\rho}\tilde{h}_{\alpha\beta}\partial_{\sigma}\tilde{h}_{\gamma\delta}&=\frac{1}{2}\tilde{h}^{\mu}_{\nu}\partial_{\mu}\tilde{h}\partial^{\nu}\tilde{h}-\frac{1}{4}\tilde{h}\partial_{\rho}\tilde{h}\partial^{\rho}\tilde{h}+\tilde{h}^{\mu}_{\nu}\partial_{\rho}\tilde{h}^{\nu}_{\mu}\partial^{\rho}\tilde{h}-\tilde{h}^{\mu}_{\nu}\partial^{\sigma}\tilde{h}^{\nu}_{\mu}\partial_{\rho}\tilde{h}^{\rho}_{\sigma}\\ &+\frac{1}{4}\tilde{h}\partial_{\rho}\tilde{h}^{\mu}_{\nu}\partial^{\rho}\tilde{h}^{\nu}_{\mu}-\tilde{h}^{\nu}_{\mu}\partial_{\nu}\tilde{h}^{\mu}_{\sigma}\partial^{\sigma}\tilde{h}-\tilde{h}^{\mu}_{\nu}\partial^{\nu}\tilde{h}\partial_{\rho}\tilde{h}^{\rho}_{\mu}+\frac{1}{2}\tilde{h}\partial_{\rho}\tilde{h}^{\rho}_{\sigma}\partial^{\sigma}\tilde{h}\\ &-\tilde{h}^{\mu}_{\nu}\partial_{\sigma}\tilde{h}^{\rho}_{\mu}\partial^{\sigma}\tilde{h}^{\nu}_{\rho}-\frac{1}{2}\tilde{h}\partial_{\mu}\tilde{h}^{\rho}_{\nu}\partial^{\nu}\tilde{h}^{\mu}_{\rho}+\tilde{h}^{\mu\nu}\partial_{\rho}\tilde{h}^{\sigma}_{\mu}\partial_{\sigma}\tilde{h}^{\rho}_{\nu}-\frac{1}{2}\tilde{h}^{\mu}_{\nu}\partial_{\mu}\tilde{h}^{\rho}_{\sigma}\partial^{\nu}\tilde{h}^{\sigma}_{\rho}\\ &+2\tilde{h}^{\mu}_{\nu}\partial^{\nu}\tilde{h}^{\sigma}_{\rho}\partial_{\sigma}\tilde{h}^{\rho}_{\mu}.\end{split}$ (18)
To obtain a manifestly covariant form, we symmetrize Eq. (17) by summing over cyclic permutations of the graviton fields $\tilde{h}_{\mu\nu}(l_{1})$, $\tilde{h}_{\alpha\beta}(l_{2})$, and $\tilde{h}_{\gamma\delta}(l_{3})$. This procedure systematically ensures the complete symmetry of the three-graviton vertex required by the underlying diffeomorphism invariance, yielding the compact expression
$\begin{split}S_{\text{GR}}=&-\frac{\kappa}{3}\int\frac{\,\mathrm{d}^{D}l_{1}}{(2\pi)^{D}}\int\frac{\,\mathrm{d}^{D}l_{2}}{(2\pi)^{D}}\int\frac{\,\mathrm{d}^{D}l_{3}}{(2\pi)^{D}}(2\pi)^{D}\delta^{D}\left(l_{1}+l_{2}+l_{3}\right)\left(U^{\mu\nu\alpha\beta\rho\gamma\delta\sigma}_{\text{GR}}l_{2\rho}l_{3\sigma}\right.\\ &\left.+U^{\alpha\beta\gamma\delta\rho\mu\nu\sigma}_{\text{GR}}l_{3\rho}l_{1\sigma}+U^{\gamma\delta\mu\nu\rho\alpha\beta\sigma}_{\text{GR}}l_{2\rho}l_{3\sigma}\right)\tilde{h}_{\mu\nu}(l_{1})\tilde{h}_{\alpha\beta}(l_{2})\tilde{h}_{\delta\gamma}(l_{3}).\end{split}$ (19)
From the above equation, we can extract the Feynman rules for the three-graviton vertex as
$\begin{split}2i\kappa V^{\mu\nu\alpha\beta\gamma\delta}_{h^{3}\text{GR}}\left(l_{1},l_{2},l_{3}\right)&=-2i\kappa\left(U^{\mu\nu\alpha\beta\rho\gamma\delta\sigma}_{\text{GR}}l_{2\rho}l_{3\sigma}+U^{\alpha\beta\gamma\delta\rho\mu\nu\sigma}_{\text{GR}}l_{3\rho}l_{1\sigma}+U^{\gamma\delta\mu\nu\rho\alpha\beta\sigma}_{\text{GR}}l_{2\rho}l_{3\sigma}\right).\end{split}$ (20)
Having established the GR contribution, we now turn to the vertex terms originating from the nonlocal sector of the action, which are expressed as
$\begin{split}S_{\text{Nonlocal}}&=\frac{2}{\kappa^{2}}\int\sqrt{-g}\,\mathrm{d}^{D}x\left(RF_{1}(\Box)R-2R_{\mu\nu}F_{1}(\Box)R^{\mu\nu}\right)\\ &=\kappa\int\,\mathrm{d}^{D}x\left(U^{\mu\nu a_{1}a_{2}abb_{1}b_{2}cd}_{1\text{Nonlocal}}\tilde{h}_{\mu\nu}\partial_{a}\partial_{b}\tilde{h}_{a_{1}a_{2}}F_{1}(\Box)\partial_{c}\partial_{d}\tilde{h}_{b_{1}b_{2}}\right.\\ &+\left.U^{\mu\nu aa_{1}a_{2}bb_{1}b_{2}cd}_{2\text{Nonlocal}}\partial_{a}\tilde{h}_{\mu\nu}\partial_{b}\tilde{h}_{a_{1}a_{2}}F_{1}(\Box)\partial_{c}\partial_{d}\tilde{h}_{b_{1}b_{2}}\right),\end{split}$ (21)
with definition
$\begin{split}&U^{\mu\nu a_{1}a_{2}abb_{1}b_{2}cd}_{1\text{Nonlocal}}\tilde{h}_{\mu\nu}\partial_{a}\partial_{b}\tilde{h}_{a_{1}a_{2}}F_{1}(\Box)\partial_{c}\partial_{d}\tilde{h}_{b_{1}b_{2}}=\tilde{h}\partial_{a}\partial_{b}\tilde{h}^{ab}F_{1}(\Box)\partial_{\mu}\partial_{\nu}\tilde{h}^{\mu\nu}-\tilde{h}\partial_{a}\partial_{b}\tilde{h}^{ab}F_{1}(\Box)\Box\tilde{h}\\ &-\tilde{h}\Box\tilde{h}F_{1}(\Box)\partial_{\mu}\partial_{\nu}\tilde{h}^{\mu\nu}+\tilde{h}\Box\tilde{h}F_{1}(\Box)\Box\tilde{h}+4\tilde{h}_{ab}\Box\tilde{h}^{ab}F_{1}(\Box)\partial_{\mu}\partial_{\nu}\tilde{h}^{\mu\nu}-8\tilde{h}_{\mu\nu}\partial^{\mu}\partial^{\rho}\tilde{h}^{\nu}_{\rho}F_{1}(\Box)\partial_{a}\partial_{b}\tilde{h}^{ab}\\ &+4\tilde{h}_{\mu\nu}\partial^{\mu}\partial^{\nu}\tilde{h}F_{1}(\Box)\partial_{a}\partial_{b}\tilde{h}^{ab}-4\tilde{h}_{\mu\nu}\Box\tilde{h}^{\mu\nu}F_{1}(\Box)\Box\tilde{h}+8\tilde{h}_{\mu\nu}\partial^{\mu}\partial^{\rho}\tilde{h}^{\nu}_{\rho}F_{1}(\Box)\Box\tilde{h}-4\tilde{h}_{\mu\nu}\partial^{\mu}\partial^{\nu}\tilde{h}F_{1}(\Box)\Box\tilde{h}\\ &-\tilde{h}\partial_{\mu}\partial_{\rho}\tilde{h}^{\rho}_{\nu}F_{1}(\Box)\partial^{\mu}\partial_{a}\tilde{h}^{a\nu}-\tilde{h}\partial_{\mu}\partial_{\rho}\tilde{h}^{\rho}_{\nu}F_{1}(\Box)\partial^{\nu}\partial_{a}\tilde{h}^{a\mu}+\tilde{h}\partial_{\mu}\partial_{\rho}\tilde{h}^{\rho}_{\nu}F_{1}(\Box)\partial^{\mu}\partial^{\nu}\tilde{h}+\tilde{h}\partial_{\mu}\partial_{\rho}\tilde{h}^{\rho}_{\nu}F_{1}(\Box)\Box\tilde{h}^{\mu\nu}\\ &+\tilde{h}\partial_{\mu}\partial_{\nu}\tilde{h}F_{1}(\Box)\partial^{\mu}\partial_{\rho}\tilde{h}^{\rho\nu}-\frac{1}{2}\tilde{h}\partial_{\mu}\partial_{\nu}\tilde{h}F_{1}(\Box)\partial^{\mu}\partial^{\nu}\tilde{h}-\frac{1}{2}\tilde{h}\partial_{\mu}\partial_{\nu}\tilde{h}F_{1}(\Box)\Box\tilde{h}^{\mu\nu}+\tilde{h}\Box\tilde{h}_{\mu\nu}F_{1}(\Box)\partial^{\mu}\partial_{\rho}\tilde{h}^{\rho\nu}\\ &-\frac{1}{2}\tilde{h}\Box\tilde{h}_{\mu\nu}F_{1}(\Box)\partial^{\mu}\partial^{\nu}\tilde{h}-\frac{1}{2}\tilde{h}\Box\tilde{h}_{\mu\nu}F_{1}(\Box)\Box\tilde{h}^{\mu\nu}+2\tilde{h}^{a\mu}\partial_{a}\partial_{\nu}\tilde{h}F_{1}(\Box)\Box\tilde{h}^{\nu}_{\mu}+2\tilde{h}^{a\mu}\partial_{a}\partial_{\nu}\tilde{h}F_{1}(\Box)\partial^{\nu}\partial_{\mu}\tilde{h}\\ &-2\tilde{h}^{a\mu}\partial_{a}\partial_{\nu}\tilde{h}F_{1}(\Box)\partial_{\mu}\partial_{\rho}\tilde{h}^{\rho\nu}+2\tilde{h}^{a\mu}\Box\tilde{h}_{a\nu}F_{1}(\Box)\Box\tilde{h}^{\nu}_{\mu}+2\tilde{h}^{a\mu}\Box\tilde{h}_{a\nu}F_{1}(\Box)\partial^{\nu}\partial_{\mu}\tilde{h}-2\tilde{h}^{a\mu}\Box\tilde{h}_{a\nu}F_{1}(\Box)\partial_{\mu}\partial_{\rho}\tilde{h}^{\rho\nu}\\ &-2\tilde{h}^{a\mu}\partial_{\nu}\partial_{\rho}\tilde{h}^{\rho}_{a}F_{1}(\Box)\Box\tilde{h}^{\nu}_{\mu}-2\tilde{h}^{a\mu}\partial_{\nu}\partial_{\rho}\tilde{h}^{\rho}_{a}F_{1}(\Box)\partial^{\nu}\partial_{\mu}\tilde{h}+2\tilde{h}^{a\mu}\partial_{\nu}\partial_{\rho}\tilde{h}^{\rho}_{a}F_{1}(\Box)\partial_{\mu}\partial_{\lambda}\tilde{h}^{\lambda\nu}-2\tilde{h}^{a\mu}\partial_{\nu}\partial_{\rho}\tilde{h}^{\rho}_{a}F_{1}(\Box)\Box\tilde{h}^{\nu}_{\mu}\\ &-2\tilde{h}^{a\mu}\partial_{\nu}\partial_{\rho}\tilde{h}^{\rho}_{a}F_{1}(\Box)\partial^{\nu}\partial_{\mu}\tilde{h}+2\tilde{h}^{a\mu}\partial_{\nu}\partial_{\rho}\tilde{h}^{\rho}_{a}F_{1}(\Box)\partial_{\mu}\partial_{\lambda}\tilde{h}^{\lambda\nu}+2\tilde{h}_{cd}\partial_{a}\partial_{\nu}\tilde{h}^{cd}F_{1}(\Box)\Box\tilde{h}^{a\nu}-2\tilde{h}^{\rho\lambda}\partial_{\rho}\partial_{a}\tilde{h}_{\lambda\nu}F_{1}(\Box)\Box\tilde{h}^{a\nu}\\ &-2\tilde{h}^{\rho\lambda}\partial_{\rho}\partial_{\nu}\tilde{h}_{\lambda a}F_{1}(\Box)\Box\tilde{h}^{a\nu}+2\tilde{h}^{\rho\lambda}\partial_{\rho}\partial_{\lambda}\tilde{h}_{a\nu}F_{1}(\Box)\Box\tilde{h}^{a\nu}+2\tilde{h}_{cd}\partial_{a}\partial_{\nu}\tilde{h}^{cd}F_{1}(\Box)\partial^{\nu}\partial^{a}\tilde{h}-2\tilde{h}^{\rho\lambda}\partial_{\rho}\partial_{a}\tilde{h}_{\lambda\nu}F_{1}(\Box)\partial^{\nu}\partial^{a}\tilde{h}-2\tilde{h}^{\rho\lambda}\partial_{\rho}\partial_{\nu}\tilde{h}_{\lambda a}F_{1}(\Box)\partial^{\nu}\partial^{a}\tilde{h}+2\tilde{h}^{\rho\lambda}\partial_{\rho}\partial_{\lambda}\tilde{h}_{a\nu}F_{1}(\Box)\partial^{\nu}\partial^{a}\tilde{h}\\ &-2\tilde{h}^{\rho\lambda}\partial_{\rho}\partial_{\nu}\tilde{h}_{\lambda a}F_{1}(\Box)\partial^{\nu}\partial^{a}\tilde{h}+2\tilde{h}^{\rho\lambda}\partial_{\rho}\partial_{\lambda}\tilde{h}_{a\nu}F_{1}(\Box)\partial^{\nu}\partial^{a}\tilde{h}-4\tilde{h}_{cd}\partial_{a}\partial_{\nu}\tilde{h}^{cd}F_{1}(\Box)\partial^{a}\partial_{\lambda}\tilde{h}^{\lambda\nu}\\ &+8\tilde{h}^{\rho\lambda}\partial_{\rho}\partial_{a}\tilde{h}_{\lambda\nu}F_{1}(\Box)\partial_{c}\partial^{(a}\tilde{h}^{\nu)c}-4\tilde{h}^{\rho\lambda}\partial_{\rho}\partial_{\lambda}\tilde{h}_{a\nu}F_{1}(\Box)\partial^{a}\partial_{c}\tilde{h}^{c\nu},\end{split}$ (22)
$\begin{split}&U^{\mu\nu aa_{1}a_{2}bb_{1}b_{2}cd}_{2\text{Nonlocal}}\partial_{a}\tilde{h}_{\mu\nu}\partial_{b}\tilde{h}_{a_{1}a_{2}}F_{1}(\Box)\partial_{c}\partial_{d}\tilde{h}_{b_{1}b_{2}}=3\partial_{\alpha}\tilde{h}_{\mu\nu}\partial^{\alpha}\tilde{h}^{\mu\nu}F_{1}(\Box)\partial_{a}\partial_{b}\tilde{h}^{ab}-4\partial_{a}\tilde{h}^{a}_{\rho}\partial_{b}\tilde{h}^{\rho b}F_{1}(\Box)\partial_{\mu}\partial_{\nu}\tilde{h}^{\mu\nu}\\ &+4\partial_{a}\tilde{h}^{a}_{\rho}\partial^{\rho}\tilde{h}F_{1}(\Box)\partial_{\mu}\partial_{\nu}\tilde{h}^{\mu\nu}-2\partial_{\mu}\tilde{h}_{\nu\alpha}\partial^{\alpha}\tilde{h}^{\mu\nu}F_{1}(\Box)\partial_{a}\partial_{b}\tilde{h}^{ab}-\partial_{\rho}\tilde{h}\partial^{\rho}\tilde{h}F_{1}(\Box)\partial_{\mu}\partial_{\nu}\tilde{h}^{\mu\nu}+\partial_{\mu}\tilde{h}\partial^{\mu}\tilde{h}F_{1}(\Box)\Box\tilde{h}\\ &-3\partial_{\alpha}\tilde{h}_{\mu\nu}\partial^{\alpha}\tilde{h}^{\mu\nu}F_{1}(\Box)\Box\tilde{h}+4\partial_{a}\tilde{h}^{a}_{\mu}\partial_{b}\tilde{h}^{\mu b}F_{1}(\Box)\Box\tilde{h}-4\partial_{a}\tilde{h}^{a}_{\mu}\partial^{\mu}\tilde{h}F_{1}(\Box)\Box\tilde{h}+2\partial_{\mu}\tilde{h}_{\nu\alpha}\partial^{\alpha}\tilde{h}^{\mu\nu}F_{1}(\Box)\Box\tilde{h}\\ &+2\partial^{\lambda}\tilde{h}\partial_{a}\tilde{h}_{\lambda\nu}F_{1}(\Box)\Box\tilde{h}^{a\nu}-\partial^{\lambda}\tilde{h}\partial_{\lambda}\tilde{h}_{a\nu}F_{1}(\Box)\Box\tilde{h}^{a\nu}-\partial_{a}\tilde{h}^{\rho\lambda}\partial_{\nu}\tilde{h}_{\rho\lambda}F_{1}(\Box)\Box\tilde{h}^{a\nu}-\partial^{\lambda}\tilde{h}^{\rho}_{a}\partial_{\rho}\tilde{h}_{\lambda\nu}F_{1}(\Box)\Box\tilde{h}^{a\nu}\\ &+\partial^{\lambda}\tilde{h}^{\rho}_{a}\partial_{\lambda}\tilde{h}_{\rho\nu}F_{1}(\Box)\Box\tilde{h}^{a\nu}+\partial^{\rho}\tilde{h}^{\lambda}_{a}\partial_{\rho}\tilde{h}_{\lambda\nu}F_{1}(\Box)\Box\tilde{h}^{a\nu}-\partial^{\rho}\tilde{h}^{\lambda}_{a}\partial_{\lambda}\tilde{h}_{\rho\nu}F_{1}(\Box)\Box\tilde{h}^{a\nu}+2\partial^{\lambda}\tilde{h}\partial_{a}\tilde{h}_{\lambda\nu}F_{1}(\Box)\partial^{\nu}\partial^{a}\tilde{h}\\ &-\partial^{\lambda}\tilde{h}\partial_{\lambda}\tilde{h}_{a\nu}F_{1}(\Box)\partial^{\nu}\partial^{a}\tilde{h}-\partial_{a}\tilde{h}^{\rho\lambda}\partial_{\nu}\tilde{h}_{\rho\lambda}F_{1}(\Box)\partial^{\nu}\partial^{a}\tilde{h}-\partial^{\lambda}\tilde{h}^{\rho}_{a}\partial_{\rho}\tilde{h}_{\lambda\nu}F_{1}(\Box)\partial^{\nu}\partial^{a}\tilde{h}+\partial^{\lambda}\tilde{h}^{\rho}_{a}\partial_{\lambda}\tilde{h}_{\rho\nu}F_{1}(\Box)\partial^{\nu}\partial^{a}\tilde{h}\\ &+\partial^{\rho}\tilde{h}^{\lambda}_{a}\partial_{\rho}\tilde{h}_{\lambda\nu}F_{1}(\Box)\partial^{\nu}\partial^{a}\tilde{h}-\partial^{\rho}\tilde{h}^{\lambda}_{a}\partial_{\lambda}\tilde{h}_{\rho\nu}F_{1}(\Box)\partial^{\nu}\partial^{a}\tilde{h}-4\partial^{\lambda}\tilde{h}\partial_{a}\tilde{h}_{\lambda\nu}F_{1}(\Box)\partial_{c}\partial^{(a}\tilde{h}^{\nu)c}+2\partial^{\lambda}\tilde{h}\partial_{\lambda}\tilde{h}_{a\nu}F_{1}(\Box)\partial^{a}\partial_{c}\tilde{h}^{c\nu}\\ &+2\partial_{a}\tilde{h}^{\rho\lambda}\partial_{\nu}\tilde{h}_{\rho\lambda}F_{1}(\Box)\partial_{c}\partial^{(a}\tilde{h}^{\nu)c}+2\partial^{\lambda}\tilde{h}^{\rho}_{a}\partial_{\rho}\tilde{h}_{\nu\lambda}F_{1}(\Box)\partial_{c}\partial^{(a}\tilde{h}^{\nu)c}-2\partial^{\lambda}\tilde{h}^{\rho}_{a}\partial_{\lambda}\tilde{h}_{\nu\rho}F_{1}(\Box)\partial_{c}\partial^{(a}\tilde{h}^{\nu)c}\\ &-2\partial^{\rho}\tilde{h}^{\lambda}_{a}\partial_{\rho}\tilde{h}_{\nu\lambda}F_{1}(\Box)\partial_{c}\partial^{(a}\tilde{h}^{\nu)c}+2\partial^{\rho}\tilde{h}^{\lambda}_{a}\partial_{\lambda}\tilde{h}_{\nu\rho}F_{1}(\Box)\partial_{c}\partial^{(a}\tilde{h}^{\nu)c}+2\partial_{a}\tilde{h}_{cd}\partial_{\nu}\tilde{h}^{cd}F_{1}(\Box)\Box\tilde{h}^{a\nu}\\ &-2\partial_{\rho}\tilde{h}^{\rho\lambda}\partial_{a}\tilde{h}_{\lambda\nu}F_{1}(\Box)\Box\tilde{h}^{a\nu}-2\partial_{\rho}\tilde{h}^{\rho\lambda}\partial_{\nu}\tilde{h}_{\lambda a}F_{1}(\Box)\Box\tilde{h}^{a\nu}+2\partial_{\rho}\tilde{h}^{\rho\lambda}\partial_{\lambda}\tilde{h}_{a\nu}F_{1}(\Box)\Box\tilde{h}^{a\nu}\\ &+2\partial_{a}\tilde{h}_{cd}\partial_{\nu}\tilde{h}^{cd}F_{1}(\Box)\partial^{\nu}\partial^{a}\tilde{h}-2\partial_{\rho}\tilde{h}^{\rho\lambda}\partial_{a}\tilde{h}_{\lambda\nu}F_{1}(\Box)\partial^{\nu}\partial^{a}\tilde{h}-2\partial_{\rho}\tilde{h}^{\rho\lambda}\partial_{\nu}\tilde{h}_{\lambda a}F_{1}(\Box)\partial^{\nu}\partial^{a}\tilde{h}\\ &+2\partial_{\rho}\tilde{h}^{\rho\lambda}\partial_{\lambda}\tilde{h}_{a\nu}F_{1}(\Box)\partial^{\nu}\partial^{a}\tilde{h}-4\partial_{a}\tilde{h}_{cd}\partial_{\nu}\tilde{h}^{cd}F_{1}(\Box)\partial_{c}\partial^{(a}\tilde{h}^{\nu)c}+4\partial_{\rho}\tilde{h}^{\rho\lambda}\partial_{a}\tilde{h}_{\lambda\nu}F_{1}(\Box)\partial_{c}\partial^{(a}\tilde{h}^{\nu)c}\\ &+4\partial_{\rho}\tilde{h}^{\rho\lambda}\partial_{\nu}\tilde{h}_{\lambda a}F_{1}(\Box)\partial_{c}\partial^{(a}\tilde{h}^{\nu)c}-4\partial_{\rho}\tilde{h}^{\rho\lambda}\partial_{\lambda}\tilde{h}_{a\nu}F_{1}(\Box)\partial_{c}\partial^{(a}\tilde{h}^{\nu)c}.\end{split}$ (23)
Similarly, the tensor $U^{\mu\nu a_{1}a_{2}abb_{1}b_{2}cd}_{1\text{Nonlocal}}$ and $U^{\mu\nu aa_{1}a_{2}bb_{1}b_{2}cd}_{2\text{Nonlocal}}$ also exhibit symmetry under the exchange of $a_{1}a_{2}ab\leftrightarrow b_{1}b_{2}cd$, $\mu\nu a\leftrightarrow a_{1}a_{2}b$, $\mu\leftrightarrow\nu$, $a_{1}\leftrightarrow a_{2}$, $a\leftrightarrow b$, $b_{1}\leftrightarrow b_{2}$ and $c\leftrightarrow d$. Note that the parentheses here indicate the symmetrization operation on the tensor, denoted as $T^{(ab)}=\frac{1}{2}\left(T^{ab}+T^{ba}\right)$. Following a similar approach as in the GR case, we perform the Fourier transform and ultimately derive the Feynman rules for the three-graviton vertex contribution from the nonlocal terms
$\begin{split}2i\kappa V^{\mu\nu a_{1}a_{2}b_{1}b_{2}}_{h^{3}\text{Nonlocal}}\left(l_{1},l_{2},l_{3}\right)&=2i\kappa\left(U^{\mu\nu a_{1}a_{2}abb_{1}b_{2}cd}_{1\text{Nonlocal}}l_{2a}l_{2b}F_{1}(-l^{2}_{3})l_{3c}l_{3d}+U^{a_{1}a_{2}b_{1}b_{2}ab\mu\nu cd}_{1\text{Nonlocal}}l_{3a}l_{3b}F_{1}(-l^{2}_{1})l_{1c}l_{1d}\right.\\ &\left.U^{b_{1}b_{2}\mu\nu aba_{1}a_{2}cd}_{1\text{Nonlocal}}l_{1a}l_{1b}F_{1}(-l^{2}_{2})l_{2c}l_{2d}+U^{\mu\nu aa_{1}a_{2}bb_{1}b_{2}cd}_{2\text{Nonlocal}}l_{1a}l_{2b}F_{1}(-l^{2}_{3})l_{3c}l_{3d}\right.\\ &\left.+U^{a_{1}a_{2}ab_{1}b_{2}b\mu\nu cd}_{2\text{Nonlocal}}l_{2a}l_{3b}F_{1}(-l^{2}_{1})l_{1c}l_{1d}+U^{b_{1}b_{2}a\mu\nu ba_{1}a_{2}cd}_{2\text{Nonlocal}}l_{3a}l_{1b}F_{1}(-l^{2}_{2})l_{2c}l_{2d}\right).\end{split}$ (24)
Eqs. (20) and (24) together contribute to the three-graviton interaction in super-renormalizable gravity, where all three gravitons are off-shell. The corresponding diagram is shown in Fig. 4.
Figure 4: The Feynman diagram for the three-graviton vertex interaction with $2i\kappa V^{\mu\nu\alpha\beta\gamma\delta}_{h^{3}}\left(p,q,k\right)$.
Appendix B Triangle Loop Integrals
We now turn to the triangle integrals governing the one-loop correction, shown in Fig. 2, beginning with their form in GR. Following the definitions in Eqs. (LABEL:28)-(LABEL:32), we first analyze the fundamental structure of these integrals by temporarily setting aside explicit vertex factors. The simplest and most illustrative case is the scalar triangle integral
$\begin{split}I^{\text{GR}}=\int\frac{\mathrm{d}^{D}l}{(2\pi)^{D}}\frac{e^{-\frac{l^{2}}{M^{2}_{\ast}}}e^{-\frac{(l+q_{\bot})^{2}}{M^{2}_{\ast}}}}{l^{2}(l+q_{\bot})^{2}[(l+k)^{2}-m^{2}+i\epsilon]},\end{split}$ (25)
where the exponential factors $e^{-l^{2}/M^{2}_{*}}$ and $e^{-(l+q_{\bot})^{2}/M^{2}_{*}}$ implement the weak nonlocality of the gravitational interaction. These form factors ensure UV convergence while preserving covariance, and become crucial in the classical limit where they regulate the short-distance behavior without introducing additional poles. The integral represents the one-loop exchange of two nonlocal graviton propagators between massive scalar propagators, which reduces to the standard GR expression in the limit $M_{*}\to\infty$.
The classical limit enables us to reduce the complexity of the massive propagator
$\begin{split}\frac{1}{(l+k)^{2}-m^{2}+i\epsilon}\approx\frac{1}{2kl+i\epsilon}=\frac{1}{2ml_{||}}-\frac{i\pi}{2m}\delta{(l_{||})}.\end{split}$ (26)
By substituting this equation into the scalar triangle integral of Eq. (25), we can disregard the first term in Eq. (26). This is because the two graviton propagators are even in $l_{||}$. In the classical limit, the scalar triangle integral simplifies to
$\begin{split}I^{\text{GR}}=-\frac{i}{4m}N_{D-1}=-\frac{i}{4m}\int\frac{{\,\mathrm{d}^{D-1}l_{\bot}}}{(2\pi)^{D-1}}\frac{e^{-\frac{l^{2}_{\bot}}{M^{2}_{\ast}}}e^{-\frac{\left(l_{\bot}+q_{\bot}\right)^{2}}{M^{2}_{\ast}}}}{l^{2}_{\bot}(l_{\bot}+q_{\bot})^{2}}.\end{split}$ (27)
The integral $N_{D-1}$ defined in Eq. (27) plays a particularly important role, as it directly corresponds to the convolution structure analyzed in the main text. In position space, this convolution reduces to simple multiplication, significantly simplifying the interpretation of the nonlocal interaction. The tensor triangle integrals $I^{\text{GR}}_{\mu}$ and $I^{\text{GR}}_{\mu\nu}$ can be treated through the same approach, with only these two tensor structures contributing to the classical gravitational potential in the GR sector.
$\left\{\begin{split}&I^{\text{GR}}_{\mu}=\int\frac{{\,\mathrm{d}^{D}l}}{(2\pi)^{D}}\frac{e^{-\frac{l^{2}}{M^{2}_{\ast}}}e^{-\frac{\left(l+q_{\bot}\right)^{2}}{M^{2}_{\ast}}}l_{\mu}}{l^{2}(l+q_{\bot})^{2}\left[(l+k)^{2}-m^{2}+i\epsilon\right]},\\ &I^{\text{GR}}_{\mu\nu}=\int\frac{{\,\mathrm{d}^{D}l}}{(2\pi)^{D}}\frac{e^{-\frac{l^{2}}{M^{2}_{\ast}}}e^{-\frac{\left(l+q_{\bot}\right)^{2}}{M^{2}_{\ast}}}l_{\mu}l_{\nu}}{l^{2}(l+q_{\bot})^{2}\left[(l+k)^{2}-m^{2}+i\epsilon\right]}.\end{split}\right$. (28)
The tensor integrals can be solved algebraically by proposing an ansatz. Let us first consider $I^{\text{GR}}_{\mu}$, which can be expressed as $Aq_{\bot\mu}+Bk_{\mu}$. The coefficients $A$ and $B$ are determined by the equations $q^{\mu}_{\bot}I_{\mu}=q^{2}_{\bot}A$ and $k^{\mu}I_{\mu}=m^{2}B$. Subsequently, we can apply relations like $2k^{\mu}l_{\mu}=(l+k)^{2}-m^{2}-l^{2}$ to simplify the numerators, reducing them to scalar integrals that no longer contain loop momenta in the numerator. It can be shown that $B=\frac{iN_{D}}{2m^{2}}$, with $N_{D}$ representing the $D$-dimensional integral
$\begin{split}N_{D}=\int\frac{{\,\mathrm{d}^{D}l_{E}}}{(2\pi)^{D}}\frac{e^{\frac{l^{2}_{E}}{M^{2}_{\ast}}}e^{\frac{\left(l_{E}+q_{\bot}\right)^{2}}{M^{2}_{\ast}}}}{l^{2}_{E}(l_{E}+q_{\bot})^{2}},\end{split}$ (29)
where the integral discussed above belong to the category of nonlocal integrals, making the conventional Wick rotation in QFT inapplicable. However, an alternative method can be employed to handle such integral Buoninfante:2022krn . Specifically, the integration variable $l_{0}$ can be defined on the imaginary axis, which is equivalent to condition $l_{0}=il_{D}$ ($l^{2}=-l^{2}_{E}$). Subsequently, the scattering amplitude calculated on the imaginary axis is mapped to the physical conclusion on the real axis.
Additionally, we can also conclude that $A=\frac{iN_{D-1}}{8m}$. Since this is a spatial integral, no above integration steps are necessary. In particular, we neglect $\int\frac{{\,\mathrm{d}^{D-1}l_{\bot}}}{(2\pi)^{D-1}}\frac{e^{-\frac{l^{2}_{\bot}}{M^{2}_{\ast}}}e^{-\frac{\left(l_{\bot}+q_{\bot}\right)^{2}}{M^{2}_{\ast}}}}{l^{2}_{\bot}}$, $\int\frac{{\,\mathrm{d}^{D-1}l_{\bot}}}{(2\pi)^{D-1}}\frac{e^{-\frac{l^{2}_{\bot}}{M^{2}_{\ast}}}e^{-\frac{\left(l_{\bot}+q_{\bot}\right)^{2}}{M^{2}_{\ast}}}}{(l_{\bot}+q_{\bot})^{2}}$ in the proof, as they do not contribute non-analytic terms in the classical limit. The final expression for $I^{\text{GR}}_{\mu}$ becomes
$\begin{split}I^{\text{GR}}_{\mu}=\frac{iN_{D-1}}{8m}q_{\bot\mu}+\frac{iN_{D}}{2m^{2}}k_{\mu}.\end{split}$ (30)
Ultimately, we apply a similar approach to derive
$\begin{split}I^{\text{GR}}_{\mu\nu}=-\frac{iq^{2}_{\bot}N_{D-1}}{16m(D-2)}\left[\left(D-1\right)\frac{q_{\bot\mu}q_{\bot\nu}}{q^{2}_{\bot}}-\eta^{\bot}_{\mu\nu}\right]-\frac{iN_{D}}{4m^{2}}\left(k_{\mu}q_{\bot\nu}+k_{\nu}q_{\bot\mu}\right).\end{split}$ (31)
Now, we focus on the integrals associated with the contributions from nonlocal terms, which are categorized into three types as defined in Eq. (LABEL:29). Using the same approach, we directly present the final results for the integrals involved.
The integral of first type
$\begin{split}&I^{(1)}_{\mu}=\frac{iN^{(1)}_{D-1}}{8m}q_{\bot\mu}+\frac{iN^{(1)}_{D}}{2m^{2}}k_{\mu},\\ &I^{(1)}_{\mu\nu}=-\frac{iq^{2}_{\bot}N^{(1)}_{D-1}}{16m(D-2)}\left[\left(D-1\right)\frac{q_{\bot\mu}q_{\bot\nu}}{q^{2}_{\bot}}-\eta^{\bot}_{\mu\nu}\right]-\frac{iN^{(1)}_{D}}{4m^{2}}\left(k_{\mu}q_{\bot\nu}+k_{\nu}q_{\bot\mu}\right),\\ &I^{(1)}_{\mu\nu\rho}=\frac{i(D+1)N^{(1)}_{D-1}}{32m(D-2)}q_{\bot\mu}q_{\bot\nu}q_{\bot\rho}-\frac{iq^{2}_{\bot}N^{(1)}_{D-1}}{32m(D-2)}\left(q_{\bot\mu}\eta^{\bot}_{\nu\rho}+q_{\bot\nu}\eta^{\bot}_{\mu\rho}+q_{\bot\rho}\eta^{\bot}_{\mu\nu}\right)\\ &-\frac{iq^{2}_{\bot}N^{(1)}_{D}}{8m^{4}(D-1)}k_{\mu}k_{\nu}k_{\rho}+\frac{iDN^{(1)}_{D}}{8m^{2}(D-1)}\left(k_{\mu}q_{\bot\nu}q_{\bot\rho}+k_{\nu}q_{\bot\mu}q_{\bot\rho}+k_{\rho}q_{\bot\mu}q_{\bot\nu}\right)\\ &-\frac{iq^{2}_{\bot}N^{(1)}_{D}}{8m^{2}(D-1)}\left(k_{\mu}\eta^{\bot}_{\nu\rho}+k_{\nu}\eta^{\bot}_{\mu\rho}+k_{\rho}\eta^{\bot}_{\mu\nu}\right),\\ &I^{(1)}_{\mu\nu\rho\sigma}=-\frac{i}{4m}\left[\frac{(D+1)(D+3)N^{(1)}_{D-1}}{16D(D-2)}q_{\bot\mu}q_{\bot\nu}q_{\bot\rho}q_{\bot\sigma}+\frac{q^{4}_{\bot}N^{(1)}_{D-1}}{16D(D-2)}\left(\eta^{\bot}_{\mu\nu}\eta^{\bot}_{\rho\sigma}+\eta^{\bot}_{\mu\rho}\eta^{\bot}_{\nu\sigma}+\eta^{\bot}_{\mu\sigma}\eta^{\bot}_{\nu\rho}\right)\right.\\ &\left.-\frac{(D+1)q^{2}_{\bot}N^{(1)}_{D-1}}{16D(D-2)}\left(q_{\bot\mu}q_{\bot\nu}\eta^{\bot}_{\rho\sigma}+q_{\bot\mu}q_{\bot\sigma}\eta^{\bot}_{\rho\nu}+q_{\bot\mu}q_{\bot\rho}\eta^{\bot}_{\nu\sigma}+q_{\bot\rho}q_{\bot\sigma}\eta^{\bot}_{\mu\nu}+q_{\bot\nu}q_{\bot\sigma}\eta^{\bot}_{\rho\mu}+q_{\bot\nu}q_{\bot\rho}\eta^{\bot}_{\mu\sigma}\right)\right]\\ &-\frac{iq^{2}_{\bot}N^{(1)}_{D}}{8m^{4}(D-1)}\left(k_{\mu}k_{\nu}k_{\rho}q_{\bot\sigma}+k_{\sigma}k_{\mu}k_{\nu}q_{\bot\rho}+k_{\rho}k_{\sigma}k_{\mu}q_{\bot\nu}+k_{\nu}k_{\rho}k_{\sigma}q_{\bot\mu}\right)\\ &-\frac{i(D+2)N^{(1)}_{D}}{16m^{2}(D-1)}\left(k_{\mu}q_{\bot\nu}q_{\bot\rho}q_{\bot\sigma}+k_{\nu}q_{\bot\mu}q_{\bot\rho}q_{\bot\sigma}+k_{\rho}q_{\bot\nu}q_{\bot\mu}q_{\bot\sigma}+k_{\sigma}q_{\bot\nu}q_{\bot\rho}q_{\bot\mu}\right)\\ &+\frac{iq^{2}_{\bot}N^{(1)}_{D-1}}{16m^{2}(D-1)}\left(k_{\mu}\eta_{\nu\rho}q_{\bot\sigma}+k_{\mu}\eta_{\sigma\nu}q_{\bot\rho}+k_{\mu}\eta_{\rho\sigma}q_{\bot\nu}+k_{\nu}\eta_{\mu\rho}q_{\bot\sigma}+k_{\nu}\eta_{\sigma\mu}q_{\bot\rho}+k_{\nu}\eta_{\sigma\rho}q_{\bot\mu}\right.\\ &\left.+k_{\rho}\eta_{\nu\mu}q_{\bot\sigma}+k_{\rho}\eta_{\sigma\mu}q_{\bot\nu}+k_{\rho}\eta_{\nu\sigma}q_{\bot\mu}+k_{\sigma}\eta_{\nu\mu}q_{\bot\rho}+k_{\sigma}\eta_{\mu\rho}q_{\bot\nu}+k_{\sigma}\eta_{\nu\rho}q_{\bot\mu}\right).\end{split}$ (32)
The integral of second type
$\begin{split}&I^{(2)}=-\frac{i}{4m}N^{(2)}_{D-1},\\ &I^{(2)}_{\mu}=\frac{iN^{(2)}_{D-1}}{8m}q_{\bot\mu}+\frac{iN^{(2)}_{D}}{2m^{2}}k_{\mu},\\ &I^{(2)}_{\mu\nu}=-\frac{iq^{2}_{\bot}N^{(2)}_{D-1}}{16m(D-2)}\left[\left(D-1\right)\frac{q_{\bot\mu}q_{\bot\nu}}{q^{2}_{\bot}}-\eta^{\bot}_{\mu\nu}\right]-\frac{iN^{(2)}_{D}}{4m^{2}}\left(k_{\mu}q_{\bot\nu}+k_{\nu}q_{\bot\mu}\right).\end{split}$ (33)
The integral of third type
[…]
In the above expression, six integrals are present, three of which contribute to the Newtonian potential. To facilitate numerical evaluation, we can represent them as follows
$\begin{split}&N^{(1)}_{D}=\int\frac{{\,\mathrm{d}^{D}l_{E}}}{(2\pi)^{D}}\frac{e^{\frac{l^{2}_{E}}{M^{2}_{\ast}}}e^{\frac{\left(l_{E}+q_{\bot}\right)^{2}}{M^{2}_{\ast}}}\left(e^{\frac{-\left(l_{E}+q_{\bot}\right)^{2}}{M^{2}_{\ast}}}-1\right)}{l^{2}_{E}(l_{E}+q_{\bot})^{4}},\\ &N^{(2)}_{D}=-\int\frac{{\,\mathrm{d}^{D}l_{E}}}{(2\pi)^{D}}\frac{e^{\frac{l^{2}_{E}}{M^{2}_{\ast}}}e^{\frac{\left(l_{E}+q_{\bot}\right)^{2}}{M^{2}_{\ast}}}\left(e^{\frac{q^{2}_{\bot}}{M^{2}_{\ast}}}-1\right)}{l^{2}_{E}q^{2}_{\bot}(l_{E}+q_{\bot})^{2}},\\ &N^{(3)}_{D}=\int\frac{{\,\mathrm{d}^{D}l_{E}}}{(2\pi)^{D}}\frac{e^{\frac{l^{2}_{E}}{M^{2}_{\ast}}}e^{\frac{\left(l_{E}+q_{\bot}\right)^{2}}{M^{2}_{\ast}}}\left(e^{\frac{-l^{2}_{E}}{M^{2}_{\ast}}}-1\right)}{l^{4}_{E}(l_{E}+q_{\bot})^{2}},\end{split}$ (35)
and
$\begin{split}&N^{(1)}_{D-1}=-\int\frac{{\,\mathrm{d}^{D-1}l_{\bot}}}{(2\pi)^{D-1}}\frac{e^{-\frac{l^{2}_{\bot}}{M^{2}_{\ast}}}e^{-\frac{\left(l_{\bot}+q_{\bot}\right)^{2}}{M^{2}_{\ast}}}\left(e^{\frac{\left(l_{\bot}+q_{\bot}\right)^{2}}{M^{2}_{\ast}}}-1\right)}{l^{2}_{\bot}(l_{\bot}+q_{\bot})^{4}},\\ &N^{(2)}_{D-1}=-\int\frac{{\,\mathrm{d}^{D-1}l_{\bot}}}{(2\pi)^{D-1}}\frac{e^{-\frac{l^{2}_{\bot}}{M^{2}_{\ast}}}e^{-\frac{\left(l_{\bot}+q_{\bot}\right)^{2}}{M^{2}_{\ast}}}\left(e^{\frac{q^{2}_{\bot}}{M^{2}_{\ast}}}-1\right)}{l^{2}_{\bot}q^{2}_{\bot}(l_{\bot}+q_{\bot})^{2}},\\ &N^{(3)}_{D-1}=-\int\frac{{\,\mathrm{d}^{D-1}l_{\bot}}}{(2\pi)^{D-1}}\frac{e^{-\frac{l^{2}_{\bot}}{M^{2}_{\ast}}}e^{-\frac{\left(l_{\bot}+q_{\bot}\right)^{2}}{M^{2}_{\ast}}}\left(e^{\frac{l^{2}_{\bot}}{M^{2}_{\ast}}}-1\right)}{l^{4}_{\bot}(l_{\bot}+q_{\bot})^{2}}.\end{split}$ (36)